# Filter Like You Test: Data-Driven Data Filtering for CLIP Pretraining

### v1

@@ 4 Experiments
Table 1: Comparison of our method to top performing methods on the DataComp medium scale filtering benchmark. DFN-FT is our best reproduction of DFN [14]. The mark * denotes results from the original paper, and $+$ indicates evaluations using our downloaded 119M candidate pool using the authors’ provided UIDs. Table 13 compares our evaluations to the results reported in the corresponding papers.
@@
DFN+NegCLIPLoss+ [46] No 0.371 0.304 0.390 0.285 0.378
HYPE+DFN+NegCLIPLoss+ [46] No 0.381 0.310 0.392 0.284 0.378
FLYT Yes 0.316 0.251 0.331 0.276 0.323
@@
M-FLYT+SCS (${\alpha}=0.15$) Yes 0.401 0.311 0.391 0.292 0.375
M-FLYT+SCS (${\alpha}=0.25$) Yes 0.395 0.309 0.393 0.310 0.378
In this section, we present the steps we took to create our top performing dataset which achieves 40.1% zero-shot ImageNet accuracy on the DataComp medium scale benchmark. We first review key details about our training and evaluation setup (Section 4.1), then discuss the challenges, design choices and results for FLYT (Section 4.2), M-FLYT (Section 4.3), and SCS (Section 4.4). Table 1 summarizes our findings and compares them to the state-of-the-art.
@@ 4.1 General Setup
We experiment mainly with the medium scale DataComp filtering track benchmark [15]. In this benchmark, the goal is to filter a 128M example candidate pool sourced from Common Crawl [^3] https://commoncrawl.org. The filtered data is then used to train a ViT-B/32 CLIP using a fixed procedure, and the resulting model undergoes a fixed evaluation on 38 downstream tasks, with performance typically reported in four main categories: ImageNet, ImageNet distribution shifts, VTAB, and retrieval, along with the overall average across all tasks.
For our experiments we were only able to download 119M out of the 128M DataComp candidate pool. To compare our filtered datasets to those published in prior work, we reapply the DataComp training and evaluation procedure on the intersection of the published datasets and our downloaded pools. As shown in Table 13, the accuracy degradation due to the missing data is between 0.1% and 1%.
@@
For both FLYT and M-FLYT, we use a CLIP ViT-B/32 model as the reference model. Unless otherwise mentioned, we use the unfiltered DataComp medium pool as the upstream dataset, and the ImageNet training set as the downstream dataset.
To enable large-scale training of FLYT, we develope a data parallelism approach, which is detailed in Appendix A.
4.2 FLYT Results
Our scoring model consists of CLIP ViT-B/32 [33] base and a ${\mathrm{FFN}_{\mathrm{GLU}}}$ [39] head. We initialize both the scoring model base and the reference model with OpenAI’s checkpoint and keep the base frozen when training. We outline key decisions here, and give detailed training parameters and ablations in Appendix C.
We evaluate three alternatives for the downstream loss function: The CLIP loss, as defined in Equation 1 ${\ell_{\text{C}}}({f}_{{{\theta}_{+}}({{w}_{1:{B}}})}({{\hat{z}}_{1:{B^{\prime}}}}))$, standard cross entropy (CE) loss, and CE with a temperature parameter (see definitions in Section C.1). Our experiments show that CLIP loss and CE with temperature demonstrate similar performance, both outperforming the standard CE loss (Table 8). For subsequent experiments, as well as the FLYT result reported in Table 1, we use CE with temperature because it better resembles the accuracy metric used to evaluate most DataComp tasks. We found it essential to use separate learned temperature parameter for the upstream and downstream loss functions when using either the CLIP or CE with temperature loss. We initialize the downstream temperature parameter ${\tau_{\mathrm{DS}}}$ to 1/0.07, as is the default when training a CLIP model from scratch, while the initial value for the upstream temperature is set to 100 as part of the pretrained reference model weights. The downstream temperature is updated during training using $\nabla_{{\tau_{\mathrm{DS}}}}{\hat{L}_{\mathrm{down}}}$.
We tried several alternatives to our base configuration. Training the full scoring model including the ViT-B/32 base lowered the performance significantly, as did freezing the reference model during training. Using weaker models to initialize either the reference model or scoring model base produced worse results. Initializing the reference model with a stronger model has negligible effect, while initializing the scoring model base with a stronger model yielded improved results (Table 8). We also tried expanding the downstream data beyond ImageNet by including training sets from 22 DataComp tasks, but observed no significant difference in performance (Table 9). As discussed in Section 5, we believe that our FLYT training stack has significant room for improvement.
@@ 4.3 M-FLYT Results
Table 2: Performance and linear weights of each input to M-FLYT, along with baseline experiments. For HYPE [20] and NegCLIPLoss [46], we replicate the scores and mixing methodology used in the original papers.
Input Linear weight ImageNet Average
@@
FLYT 0.65 0.316 0.323
@@
Sum 0.324 0.339
Standardized Sum 0.344 0.361
For M-FLYT, we train a linear mixing model using input scores that are standardized to zero mean and unit standard deviation. The standardization both improves performance and provides interpretable weights for each input method. For M-FLYT we empirically found that initializing the reference model with a CLIP model trained using DataComp’s medium scale configuration on their unfiltered dataset and using a downstream CE loss without temperature produced better results: see Section C.6
@@
We compare M-FLYT to its individual input components and to two baseline approaches. The first baseline uses the sum of all input scores, while the second sums the scores after standardization to zero mean and unit standard deviation. Table 2 demonstrate that M-FLYT outperforms all individual scores as well as both baseline aggregations. Notably, while FLYT has modest performance as a standalone method, it receives the second-highest weight, surpassed only by DFN-FT, indicating that FLYT picks up scoring signal that are complementary to those found in prior work. Table 11 demonstrates that excluding any individual input score from M-FLYT’s training leads to performance degradation, suggesting that the method effectively leverages multiple signals, and could potentially achieve stronger performance given more (or better) input scores.
@@ 4.4 SCS Results
Figure 3: Comparison of sampling strategies (SCS, HCS, standard threshold filtering, and No Cap) showing their effect on ImageNet accuracy (top) and average accuracy (bottom). SCS was tested with ${\alpha}$ values from 0.1 to 0.6, HCS with ${\beta}$ values from 5 to 25 and standard threshold with values from 5% to 25%
For SCS, we always sample in batches of ${G}=100K$, and experiment with a repetition penalty ${\alpha}$ between 0.1 and 0.6. When preparing the filtered dataset, we use a shuffle buffer of size 1.5M in the DataComp resharder to prevents batches from containing many copies of the same example.
@@
When using a repetition penalty ${\alpha}$ of 0.15, which yields 25.67M unique examples, we achieve the highest ImageNet accuracy. However, increasing ${\alpha}$ to 0.25 generates more unique examples (29.43M) and increases average accuracy across all 38 downstream tasks. This pattern, where example diversity improves average accuracy but reduces ImageNet accuracy, also appears in threshold filtering: selecting the top 10% of examples (11.93M) performs better on ImageNet compared to selecting the top 20% (23.86M), but results in lower average accuracy across tasks.
@@ 4.5 Small Scale Results
Table 3: Comparing our methods to top performing methods on the DataComp small scale filtering benchmark. * denotes results we got from the DataComp leaderboard that did not provide a paper to cite.
@@
We evaluate our method on the small scale DataComp benchmark. We use the mixing model trained by M-FLYT described in Section 4.3, and sample using SCS with a repetition penalty ${\alpha}=0.5$ and a batch size ${G}=10K$. Table 3 shows that our method achieves 8% zero-shot ImageNet zero-shot accuracy and 18.5% average accuracy, outperforming the top methods in the DataComp leaderboard on both categories. These results suggest that M-FLYT and SCS can be effectively used across different computational scales. Figure 4 shows the effect of different ${\alpha}$ values
@@ 5 Limitations and Future Work
Another notable limitation is that FLYT training on raw inputs did not succeed, and on fixed embeddings did not surpass state-of-the-art methods (though it did contribute to M-FLYT). We believe this is likely a deficiency in our optimization stack, solvable by more careful tuning of learning rate (schedules) for individual components and regularization techniques such as dropout. In similar vein, we believe that more careful parameter tuning should allow us to leverage downstream data that is more diverse than ImageNet.
@@
A fourth direction for future work is closing the data selection / pretraining loop by iteratively re-training the scoring model to select the next data to pretrain our model on. FLYT is particularly well-suited to this form of active learning, as it allows us to use the pretraining model as the reference model at each scoring iteration, ensuring that each batch we select is most useful for the model at its current state.
@@ Appendix A Implementing data-parallel distribution
$\frac{\partial}{\partial{\phi}}{{w}_{1:{B}}}=\frac{\partial}{\partial{\phi}}{\mathrm{softmax}}({{q}_{\phi}}({{z}_{1:{B}}}))$
@@ Appendix B M-FLYT Input Details
    FLYT. We use our best performing FLYT scoring model, as detailed in Section 4.2.
@@ Appendix C Experimental Setup
C.1 Downstream loss
where ${\tau_{\mathrm{DS}}}$ is a learnable temperature parameter we only use in the CE+temperature variation for the downstream data. Using the standard CE loss, ${\tau_{\mathrm{DS}}}$ is set to 1. Using this notation, we define the downstream CLIP loss mentioned in Section 4.2 with a separate temperature parameter ${\tau_{\mathrm{DS}}}$ as
C.2 Experiment configurations
Table 6: Detailed hyper parameters of the different configurations used for training and ablations of FLYT and M-FLYT. * Initialization “DataComp medium” means we train a model from scratch on the unfiltered DataComp medium scale dataset using the DataComp medium scale training configuration.
Parameter FLYT-Base FLYT-22 FLYT-LR M-FLYT-Base
@@ C.2 Experiment configurations
Scoring model base init OpenAI OpenAI OpenAI None
@@
Trainable scoring parameters only head only head only head all
@@
When training FLYT and M-FLYT, we optimize both the scoring model, the mixing model and the reference models using AdamW [24] with a weight decay of 0.2, $(\beta_{1},\beta_{2})=(0.9,0.98)$, and a cosine scheduler [23]. Due to computational constraints, some experiments compare against model configurations that differ from those listed in Section 4. We list their full hyper parameters in Table 6 and describe their main differences here. We refer to the configurations from Section 4.2, and Section 4.3 as FLYT-Base and M-FLYT-Base respectively. FLYT-22 maintains the same parameters as FLYT-Base, with three key differences: it employs the standard CLIP loss as the downstream loss, uses a downstream batch size ${B^{\prime}}$ of 768, and trains with 22 downstream tasks. FLYT-LR uses the same configuration as FLYT-Base, except it employs the CE loss without temperature and uses a downstream batch size ${B^{\prime}}$ of 1536; this configuration was used exclusively for learning rate experiments.
C.3 FLYT Learning Rate Tuning
@@
Table 7 compares different scoring model and reference model learning rates using the FLYT-LR configuration. We notice that FLYT is not very sensitive to the scoring model LR (with reference LR fixed at 1e-4) and select 1e-3 based on slightly better results. FLYT is somewhat sensitive to reference model LR, and with scoring model LR fixed at 1e-3, 5e-5 gives the best results. These values are used for subsequent experiments.
@@
Table 7: Learning rate tuning using the FLYT-LR configuration (Table 6). We first optimize the scoring model LR with reference model LR fixed at 1e-4, then optimize reference model LR with scoring LR fixed at 1e-3.
@@
C.4 FLYT Ablations
Table 8 presents experiments using FLYT-Base. It shows a comparison between downstream losses, where using CE with a temperature parameter is comparable to the CLIP downstream loss, both outperforming the standard CE loss. It compares reference model and scoring model base initializations, where initializing with stronger models outperform weaker ones. We evaluate four initialization methods in ascending order of performance: Random initialization, “DataComp medium”, where we train a model from scratch on the unfiltered DataComp medium scale dataset using the DataComp medium scale training configuration, OpenAI checkpoint, and “DataComp-13B”, which is a model trained by Gadre et al. [15] on the DC-1B dataset for 13B samples seen. Table 8 also shows that training the reference model is much better than freezing it, but freezing the scoring model base is required for the training to succeed. Our ablation studies show that performance remains stable with downstream batch sizes ranging from 1024 to 4096. Similarly, performance is consistent when training for 1250 to 12500 steps, suggesting that FLYT could be optimized using less compute.
Table 8: Experiments using the FLYT-Base configuration (Table 6). * Initialization “Trained on DC-1B” means using the ViT-B-32 model trained by Gadre et al. [15] on DC-1B for 13B samples seen (Hugging Face model card).

### v2

@@ 4 Experiments
Table 1: Comparison of our method to top performing methods on the DataComp medium scale filtering benchmark. DFN-FT is our best reproduction of DFN [16]. * denotes results from the original paper, and $+$ indicates evaluations using our downloaded 119M candidate pool using the authors’ provided UIDs. Table 14 compares our evaluations to the results reported in the corresponding papers.
@@
DFN+s-CLIPLoss+ [49] No 0.371 0.304 0.390 0.285 0.378
HYPE+DFN+s-CLIPLoss+ [49] No 0.381 0.310 0.392 0.284 0.378
@@
M-FLYT+SCS Yes 0.401 0.311 0.396 0.292 0.377
In this section, we present the steps we took to create our top performing dataset which achieves 40.1% zero-shot ImageNet accuracy on the DataComp medium scale benchmark and 37.7% average accuracy across all DataComp evaluation tasks. We first review key details about our training and evaluation setup (Section 4.1), then discuss the challenges, design choices and results for M-FLYT (Section 4.2) and SCS (Section 4.3). Table 1 summarizes our findings and compares them to the state-of-the-art. In Section 4.4 we apply our algorithm to the DataComp small scale benchmark, demonstrating that our approach works effectively across different data scales.
@@ 4.1 General setup
We experiment mainly with the medium scale DataComp filtering track benchmark [17]. In this benchmark, the goal is to filter a 128M example candidate pool sourced from Common Crawl [1]. The filtered data is then used to train a ViT-B/32 CLIP using a fixed procedure, and the resulting model undergoes a fixed evaluation on 38 downstream tasks, with performance typically reported in four main categories: ImageNet, ImageNet distribution shifts, VTAB, and retrieval, along with the overall average across all tasks.
For our experiments we were only able to download 119M out of the 128M DataComp candidate pool. To compare our filtered datasets to those published in prior work, we reapply the DataComp training and evaluation procedure on the intersection of the published datasets and our downloaded pools. Table 14 shows that the accuracy degradation due to the missing data is between 0.1% and 1%.
@@
For both M-FLYT (Section 4.2) and E-FLYT (Section C.7), we use a CLIP ViT-B/32 model as the reference model. Unless otherwise mentioned, we use the unfiltered DataComp medium pool as the upstream dataset, and the ImageNet training set as the downstream dataset.
To enable large-scale training of FLYT, we develope a data parallelism approach, which is detailed in Appendix A. The computational costs are outlined in Section C.2.
@@ 4.2 M-FLYT results
Table 2: Left: Performance and linear weight of each input to M-FLYT. For HYPE [22] and s-CLIPLoss [49], we replicate the scores and mixing methodology used in the original papers. Right: Baseline aggregation experiments as described in Section 4.2.
Input Weight ImageNet Average
@@
E-FLYT 0.65 0.316 0.323
Baseline ImageNet Average
Sum 0.327 0.331
Standardized Sum 0.348 0.360
IN-weighted ${r}=2$ 0.350 0.364
IN-weighted ${r}=4$ 0.354 0.359
IN-weighted ${r}=8$ 0.353 0.365
IN-weighted ${r}=16$ 0.352 0.362
@@
We train a linear mixing model using input scores that are standardized to zero mean and unit standard deviation. The standardization both improves performance and provides interpretable weights for each input method. We empirically found that initializing the reference model with a CLIP model trained using DataComp’s medium scale configuration on their unfiltered dataset and using standard cross entropy (CE) as the downstream loss produced better results; see Section C.3
@@
We compare M-FLYT to its individual input components and to several baseline approaches. These baselines include a simple sum of all input scores and a sum of scores after standardization to zero mean and unit standard deviation. Finding that standardization significantly improves performance, we further explore weighted standardized sums where inputs are weighted by their normalized ImageNet scores adjusted to achieve a specific max-to-min weight ratio ${r}$, for details see Section C.5.
Table 2 demonstrate that M-FLYT outperforms all individual scores as well as all baseline aggregations. Table 8 demonstrates that excluding any individual input score from M-FLYT’s training leads to performance degradation, suggesting that the method effectively leverages multiple signals, and could potentially achieve stronger performance given more (or better) input scores.
@@ 4.3 SCS results
Figure 3: Comparison of sampling strategies (SCS, HCS, threshold filtering, and No Cap) showing their effect on ImageNet accuracy (left) and average accuracy (right). SCS was tested with ${\alpha}$ values from 0.1 to 0.6, HCS with ${\beta}$ values from 5 to 25 and standard threshold with values from 5% to 25%
For SCS, we experiment with a repetition penalty ${\alpha}$ between 0.1 and 0.6, and sample in batches of ${G}=100K$. We choose ${G}$ to balance efficient sampling while ensuring the resulting dataset closely approximates what we would get using ${G}=1$. Section C.10 shows that the exact value of ${G}$ has minimal impact on performance. When preparing the filtered dataset, we use a 1.5M shuffle buffer in the DataComp resharder to prevent batches from containing many copies of the same example.
@@
Figure 3 also shows a pattern where for threshold filtering, example diversity improves average accuracy but reduces ImageNet accuracy: selecting the top 10% of examples (11.93M) performs better on ImageNet compared to selecting the top 20% (23.86M), but results in lower average accuracy across tasks. This trade-off is not as apparent with SCS.
@@ 4.4 Small scale results
Table 3: Comparing our methods to top performing methods on the DataComp small scale filtering benchmark. * denotes results from the DataComp leaderboard that did not provide a paper to cite.
@@
We evaluate our method on the small scale DataComp benchmark. We use the mixing model trained by M-FLYT described in Section 4.2, and sample using SCS with a repetition penalty ${\alpha}=0.5$ and a batch size ${G}=10K$. Table 3 shows that our method achieves 8% zero-shot ImageNet zero-shot accuracy and 18.5% average accuracy, outperforming previous work on both categories. These results suggest that M-FLYT and SCS can be effectively used across different computational scales. Figure 4 shows the effect of different ${\alpha}$ values.
@@ 5 Limitations and future work
Another notable limitation is that E-FLYT underperforms, though its scores did contribute to M-FLYT. To surpass state-of-the-art methods we had to leverage existing filtering methods through M-FLYT rather than relying on a general feature extractor. We believe this is likely a deficiency in our optimization stack, solvable by more careful tuning of learning rate (schedules) for individual components and regularization techniques such as dropout. In similar vein, we believe that more careful parameter tuning should allow us to leverage downstream data that is more diverse than ImageNet, though these hypotheses require empirical validation.
@@
A fourth direction for future work is closing the data selection / pretraining loop by iteratively re-training the scoring model to select the next data to pretrain our model on. FLYT is particularly well-suited to this form of active learning, as it allows us to use the pretraining model as the reference model at each scoring iteration, ensuring each selected batch is optimal for the model’s current state.
@@ Appendix A Implementing data-parallel distribution
$\frac{\partial}{\partial{\phi}}{{w}_{1:{B}}}=\frac{\partial}{\partial{\phi}}{\mathrm{softmax}}({{q}_{\phi}}({{\zeta}_{1:{B}}}))$
@@ Appendix B M-FLYT input details
    E-FLYT. We use our E-FLYT-Base scoring model, as detailed in Section C.7.
@@ Appendix C Experimental setup
C.1 Experiment configurations
Table 6: Detailed hyper parameters of the different configurations used for training and ablations of M-FLYT and E-FLYT. * Initialization “DataComp medium” means we train a model from scratch on the unfiltered DataComp medium scale dataset using the DataComp medium scale training configuration.
Parameter E-FLYT-Base E-FLYT-22 E-FLYT-LR M-FLYT-Base
@@ C.1 Experiment configurations
Feature extractor init OpenAI OpenAI OpenAI None
@@
Trainable scoring parameters all all all all
@@
When training M-FLYT and E-FLYT, we optimize the mixing model, the embeddings scoring model, and the reference models using AdamW [26] with a weight decay of 0.2, $(\beta_{1},\beta_{2})=(0.9,0.98)$, and a cosine scheduler [25]. Due to computational constraints, some experiments compare against different model configurations. We list their full hyper parameters in Table 6 and describe their main differences here. We refer to the configuration from Section 4.2 as M-FLYT-Base, and to the E-FLYT configuration used in the experiments described in Section C.7 as E-FLYT-Base. E-FLYT-22 maintains the same parameters as E-FLYT-Base, with three key differences: it employs the standard CLIP loss as the downstream loss, uses a downstream batch size ${B^{\prime}}$ of 768, and trains with 22 downstream tasks. E-FLYT-LR uses the same configuration as E-FLYT-Base, except it employs the CE loss without temperature and uses a downstream batch size ${B^{\prime}}$ of 1536; this configuration was used exclusively for learning rate experiments.
C.2 Computational cost
In terms of FLOPs, M-FLYT requires slightly more than 2×2$\times$2 × the FLOPs of regular CLIP model training per example, while E-FLYT requires slightly more than 3×3$\times$3 ×. Since we only train our models for 20M examples (compared to DataComp’s 128M), M-FLYT uses approximately 3×3$\times$3 × fewer FLOPs than training a DataComp medium scale model, and E-FLYT uses approximately 2×2$\times$2 × fewer.
With our implementation, M-FLYT takes 40 A100 hours to train, and E-FLYT takes 45 A100 hours, roughly equivalent to running a DataComp medium scale experiment. For SCS with a batch size ${G}=100K$, it takes about 100 minutes to sample 128M examples using a naive Python implementation on a single CPU. For larger scales (DataComp large or Xlarge), the filtering computation becomes negligible relative to the overall training cost.
Comparing to other filtering methods:
  • Fang et al. [16] reported (in Table 2) training their ViT-B/32 DFN for 5.12B samples seen, which requires approximately 80-120x more FLOPs than our method
  • Kim et al. [22] trained a ViT-L/14 CLIP model on 8 V100 GPUs for 256M samples, taking 488 GPU hours (61 hours $\times$ 8 GPUs)
  • Wang et al. [49], which doesn’t require additional model training beyond the pretrained OpenAI CLIP model, reported 5 hours of preprocessing on an L40 machine
C.3 M-FLYT ablations
Table 7 shows experiments using M-FLYT. The standard CE loss slightly outperforms CE with temperature (Section C.8), and reference model initialized with “DataComp medium” (explained in Section C.7) outperforms OpenAI initialization. We also find that input standardization greatly increases performance.
Table 7: Experiments using the M-FLYT-Base configuration (Table 6).
Experiment Baseline Variation ImageNet Average
M-FLYT 0.359 0.371
Downstream loss CE CE + temperature 0.352 0.366
OpenAI 0.342 0.355
Input standardization Yes No 0.323 0.337
Inputs ImageNet ImageNet dist. shifts VTAB Retrieval Average
All Inputs 0.359 0.294 0.383 0.310 0.371
[…]
No E-FLYT 0.352 (0.007) 0.292 (0.002) 0.376 (0.007) 0.309 (0.001) 0.366 (0.005)
C.4 Comparing M-FLYT to aggregation methods in prior work
Table 9 compares M-FLYT against the original mixing strategies of HYPE and s-CLIPLoss. For HYPE’s components, M-FLYT outperforms the simple summation approach on both ImageNet (+1.3%) and average (+0.6%) metrics. When mixing s-CLIPLoss scores, M-FLYT performs similarly to the original two-stage filtering method, showing marginal improvement on ImageNet (0.3%) but slightly lower average performance (0.6%).
Table 9: Comparing M-FLYT to the original mixing method of HYPE [22] and s-CLIPLoss [49]. HYPE combines scores by summing their three components (${-d_{L}}+{\epsilon_{i}}+{\epsilon_{t}}$), while s-CLIPLoss applies a two-stage filtering process: selecting the top 30% based on s-CLIPLoss scores, then selecting the top 20% according to their NormSim scores from this filtered subset.
Input method Mixing method ImageNet Average
HYPE Original (sum scores) 0.309 0.341
M-FLYT 0.322 0.347
s-CLIPLoss and NormSim Original (two stage) 0.331 0.363
M-FLYT 0.334 0.357
C.5 M-FLYT baseline - ImageNet-weighted standardized sum
Here we explain the ImageNet-weighted standardized sum baseline we compare against in Section 4.2. For $K$ ImageNet accuracies $\mathrm{IN}_{1:K}$ corresponding to $K$ scoring methods, we first standardize the per-example scores produced by these methods individually across all training examples to have zero mean and unit standard deviation. We denote by $s^{j}_{i}$ the standardized score that method $i\in{1,...,K}$ gave example ${z}_{j}$.
We then normalized the ImageNet accuracies $\mathrm{IN}_{1:K}$ between zero and one, and add a scalar to achieve our desired ratio ${r}$:
$\hat{w}_{i}$ $=\frac{\mathrm{IN}_{i}-\mathrm{min}(\mathrm{IN}_{1:K})}{\mathrm{max}(\mathrm{IN}_{1:K})-\mathrm{min}(\mathrm{IN}_{1:K})},$
$w_{i}$ $=\hat{w}_{i}+\frac{1}{{r}-1},$
where $\mathrm{min}(\mathrm{IN}_{1:K})$ is the lowest ImageNet score, and $\mathrm{max}(\mathrm{IN}_{1:K})$ is the highest. This ensures that the ratio between the maximum and minimum weights equals ${r}$:
${r}=\frac{\mathrm{max}(w_{1:K})}{\mathrm{min}(w_{1:K})}.$
The final score for example ${z}_{j}$ is then $\sum_{i=1}^{K}w_{i}s^{j}_{i}$.
In Table 2 we experiment with max-to-min ratios ${r}\in\{2,4,8,16\}$.
C.6 E-FLYT learning rate tuning
@@
Table 10 compares different scoring model and reference model learning rates using the E-FLYT-LR configuration. We notice that E-FLYT is not very sensitive to the scoring model LR (with reference LR fixed at 1e-4) and select 1e-3 based on slightly better results. E-FLYT is somewhat sensitive to reference model LR, and with scoring model LR fixed at 1e-3, 5e-5 gives the best results. These values are used for subsequent experiments.
@@
Table 10: Learning rate tuning using the E-FLYT-LR configuration (Table 6). We first optimize the scoring model LR with reference model LR fixed at 1e-4, then optimize reference model LR with scoring LR fixed at 1e-3.
@@
C.7 E-FLYT results
For E-FLYT, we use a CLIP ViT-B/32 [35] as the feature extractor and an ${\mathrm{FFN}_{\mathrm{GLU}}}$ [42] scoring model. We initialize both the feature extractor and the reference model with OpenAI’s checkpoint.
In Table 11 we show experiments using E-FLYT-Base. We evaluate three alternatives for the downstream loss function: The CLIP loss, as defined in Equation 1 ${\ell_{\text{C}}}({f}_{{{\theta}_{+}}({{w}_{1:{B}}})}({{\hat{z}}_{1:{B^{\prime}}}}))$, standard cross entropy (CE) loss, and CE with a temperature parameter (see definitions in Section C.8). Our experiments show that CLIP loss and CE with temperature demonstrate similar performance, both outperforming the standard CE loss. For subsequent experiments, we use CE with temperature because it better resembles the accuracy metric used to evaluate most DataComp tasks. We found it essential to use separate learned temperature parameter for the upstream and downstream loss functions when using either the CLIP or CE with temperature loss. We initialize the downstream temperature parameter ${\tau_{\mathrm{DS}}}$ to 1/0.07, as is the default when training a CLIP model from scratch, while the initial value for the upstream temperature is set to 100 as part of the pretrained reference model weights. The downstream temperature is updated during training using $\nabla_{{\tau_{\mathrm{DS}}}}{\hat{L}_{\mathrm{down}}}$.
Table 11 also compares reference model and feature extractor initializations, where initializing with stronger models outperform weaker ones. We evaluate four initializations in ascending order of performance: Random initialization, “DataComp medium”, where we train a model from scratch on the unfiltered DataComp medium scale dataset using the DataComp medium scale training configuration, OpenAI checkpoint, and “DataComp-13B”, which is a model trained by Gadre et al. [17] on the DC-1B dataset for 13B samples seen. Further results in the table show that training the reference model is much better than freezing it, and that fine-tuning the feature extractor prevents the training from succeeding. Our ablation studies show that performance remains stable with downstream batch sizes ranging from 1024 to 4096. Similarly, performance is consistent when training for 1250 to 12500 steps, suggesting that E-FLYT could be optimized using less compute.
As discussed in Section 5, we believe that our E-FLYT training stack has significant room for improvement.
Table 11: Experiments using the E-FLYT-Base configuration (Table 6). * Initialization “Trained on DC-1B” means using the ViT-B-32 model trained by Gadre et al. [17] on DC-1B for 13B samples seen (Hugging Face model card).